One fifth of 12 more than a certain number equals one-forth of 2 more than the number. Find the number
Found 2 solutions by stanbon, ptaylor:
Answer by stanbon(75887)

One fifth of 12 more than a certain number equals one-forth of 2 more than the number. Find the number
----------------------------
Let the number be "x".
EQUATION:
(1/5)(x+12) = (1/4)(x+2)
Multiply both sides by 20 to get rid of the fractions:
4(x+12) = 5(x+2)
4x+48 = 5x+10
x = 38
================
Cheers,
Stan H.

Answer by ptaylor(2198)

Let x=the number
x+12=12 more than the number
(1/5)(x+12)=one-fifth of 12 more than the number--------------a
x+2=two more than the number
(1/4)(x+2)=one fourth of two more than the number-----------------------b
Now we are told that a=b, so:
(1/5)(x+12)=(1/4)(x+2) multiply both sides by 20, the LCM
4(x+12)=5(x+2) get rid of parens
4x+48=5x+10 subtract 5x and also 48 from both sides
4x-5x+48-48=5x-5x+10-48 collect like terms
-x=-38 multiply each side by -1
x=38----the number
CK
(1/5)(50)=(1/4)(40)
10=10
Hope this helps---ptaylor